OpenStudy (anonymous):

Does anyone know the train track equation you use in chemistry? I'm trying to convert 50 Feet per 3.79 seconds to miles per hour. I just don't understand PLEASE HELP!!

OpenStudy (hlilly2413):

Try this: \[\frac{ 50 ft. }{ 3.79s }\times \left( \frac{ 1 mile }{ 5280 ft } \right) \times \left( \frac{ 60 seconds }{ 1 \min } \right) \times \left( \frac{ 60 mins }{ 1 hr } \right)\] As you go, your units should cancel (ft and ft. We use 5,280 feet because that's how many feet are in a mile. Since that is what you are converting to, you need to know that. Then you are converting seconds to minutes so your 60 seconds should cancel out the seconds next to the 3.79...Your minutes will cancel out and the units you should have left over are miles/hour.